OpenStudy (anonymous):

Solve using the addition method 2x -3y=8 3x+2y=-1

OpenStudy (anonymous):

multiply one equation by 2, and the other by three and then add or subtract the terms that have equal coefficients (you will be left with one variable, then solve for that variable and plug it into one of the original equations to get the other variable

OpenStudy (anonymous):

@killabarbie that doesnt help you any because it still has two different variables in it

OpenStudy (anonymous):

Multiply the first line by 2 multiply the second line by 3 4x -6y = 16 9x +6y = -3 The y's cancel out so you're left with 11x = 13 x = 13/11 then you can sub x into either equation and get y = -25/11

OpenStudy (anonymous):

Using matrices is long and time consuming, so dont use them.

OpenStudy (anonymous):

i couldve sworn i saw a big matrix displayed on the screen by lagrangeson

OpenStudy (anonymous):

anyone else see it? it isnt there anymore...

OpenStudy (anonymous):

Matricies is much easier actually in my opinion. also i made an error in my first calculation..

OpenStudy (anonymous):

bt you have to draw them, and then, plug the numbers in... i just dont like them because you draw out the matrix

OpenStudy (anonymous):

Trying to remember how to use matrices. Mind refreshing my memory?

OpenStudy (anonymous):

sorry ive been doing some exam studying, and I worked this equation out on paper, and the x = 1 y = -2

OpenStudy (anonymous):

Mridrik, erm I have a personal way of handling matricies that on plain paper looks longer than substitution or elimination, but with this method, u can mentally figure it out much quicker. I'll try and write an outline of it for you.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Alright so I would first set it up as a matrix. [2 -3 8] [3 2 -1] Then i would use the x values and split them to be infront of the ANS and I just keep the y - ANS format the same so i get [8 2] [-1 3] for x ______________ [-3 8] [2 -1] for y Now you divide it by the original X and Y's [2 -3] [3 2] And u multiply it all out and get your answers for x and y. so essentially its [8 2] [-1 3] DIVIDE BY [2 -3] [3 2] = y and [-3 8] [2 -1] DIVIDE BY [2 -3] [3 2] = x